### v2

# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

Abstract
Errors associated with qubit loss constitute an important source of noise in many quantum hardware systems, particularly in neutral atom quantum computers. We develop a theoretical framework to handle these errors in logical algorithms, incorporating decoding techniques and circuit-level optimizations. Focusing on experimentally-motivated error models, we introduce a delayed-erasure decoder which leverages information from state-selective readout to accurately correct loss errors, even when their precise locations are unknown. Our decoding technique is compatible with a wide range of quantum error correction codes and general logical circuits. Using this decoder, we identify strategies for detecting and correcting atom loss based on the logical circuit structure. For deep circuits with a large number of gate layers prior to logical measurements, we explore methods to integrate loss detection into syndrome extraction with minimal overhead, identifying optimal strategies depending on the qubit loss fraction in the noise. In contrast, many algorithmic subroutines involve frequent gate teleportation, shortening the circuit depth before logical measurement and naturally replacing qubits without additional overhead. We simulate such a teleportation-based algorithm, involving a toy model for small-angle synthesis and find a significant improvement in logical error rates as the loss fraction increases, with loss handled solely through teleportation. These results provide a path forward for advancing large-scale fault tolerant quantum computation in systems with loss errors.
1 Introduction
Quantum error correction (QEC) is believed to be essential for realizing large-scale quantum computation, as it enables suppression of errors [1, 2, 3]. However, its practical implementation remains challenging due to its substantial resource overhead. Recent experimental advancements have demonstrated remarkable progress in implementing quantum operations across multiple logical qubits [4, 5, 6, 7, 8], and operating below error thresholds [9]. These advances make it clear that practical QEC performance can be substantially improved by tailoring the error correction strategy to the particular experimental error model [10, 9, 11], choice of logic gates [4], and the structure of the algorithm itself [12, 13].
In particular, qubit loss and leakage from computational subspace are dominant noise sources in many hardware systems, and can have a dramatic effect on the practical performance of quantum processors [14, 15]. In the absence of loss correction, all qubits within a QEC code can eventually disappear, destroying the encoded quantum information. Conversely, directly detecting loss provides valuable information about the error location, in contrast with Pauli errors, which are indirectly inferred from syndrome information. Recent work has shown that the rich information provided by so-called erasure detection can in fact improve QEC performance substantially [16, 17, 18, 19, 10, 11, 20, 21, 22, 23, 24, 25, 26]. Similarly, noise bias present in many platforms [14, 4, 9] can also be leveraged to improve performance [27, 28, 29, 30]. These studies motivate developing strategies that incorporate loss detection and harness bias, to improve practical QEC performance.
In this Article, we explore the role of loss and biased Pauli errors in error-corrected circuits, with a focus on neutral atom quantum computers. We leverage loss detection upon measurement and design QEC techniques to benefit from loss with minimal experimental overhead. In particular, since the detection and correction of erasure is delayed for several gate operations, we develop a delayed-erasure decoder to accurately interpret logical measurement results from the measured syndromes and loss detections, despite uncertainty in the exact location of the loss error.
Using this delayed-erasure decoder, we investigate the impact of loss errors in logical circuits. We demonstrate that the algorithmic structure significantly influences the optimal strategy to detect and correct loss, as summarized in Fig. 1(a). Concretely, for high-depth circuits prior to logical measurement, we develop hardware-efficient methods to detect and replace lost qubits during syndrome extraction (SE) with minimal additional overhead. We perform circuit-level simulations to compare circuit-based [3], measurement-based [31], and Steane-based error correction [32] approaches which incorporate loss detection upon measurement, alongside other state-of-the-art approaches such as erasure conversion techniques [10, 17]. We observe that QEC performance can be substantially improved by optimizing circuit design in cases where atom loss constitutes a substantial fraction of the error budget. We then study how loss-detecting SE rounds can be optimally interleaved between transversal gates in multi-qubit deep logical Clifford circuits.
We next explore an important class of circuits with minimal operations before logical measurement, which is associated with key algorithmic subroutines such as magic state distillation [33], quantum arithmetic [34], and small-angle synthesis [35, 36]. In such circuits, gate teleportation naturally detects and replaces lost qubits without additional overhead. As a result, we find that loss-detecting SE is not needed in this setting. Our numerical simulations of a teleportation-based algorithm indicate that as the frequency of gate teleportation increases, the performance is getting closer to that of erasure conversion. These results provide a comparative analysis of experimental solutions for leveraging loss and highlight that algorithmic structure plays a central role in how these factors impact performance.
Before proceeding, we note that loss-to-erasure conversion has been explored at the memory level, using mid-circuit measurement in alkaline-earth-like atomic systems [17, 10, 16, 11] and superconducting qubits [19, 21, 20, 15, 37, 38, 39], as well as using leakage-reduction units with extra qubits and gates [25, 40, 41, 42]. Since the loss must be addressed in all systems and erasure conversion may not be always directly accessible after each gate operation (or, alternatively, may result in additional time and errors), we focus on the cases involving loss-resolving readout as a part of the natural QEC cycles. Most recently, atom loss-resolving readout was demonstrated experimentally and its implications for logical qubit performance were explored in Refs. [22, 24, 23, 8]. We extend these previous studies, which focus on the performance of individual logical qubits, to analyze atom loss effects at all levels ranging from decoding and QEC architecture to logical algorithms. We systematically compare different loss detection and correction methods, including their frequencies and associated space-time overheads. By focusing on logical algorithms we provide new insights into how circuit structure influences the algorithmic performance and develop new, optimized strategies for circumventing atom loss.
Our paper is organized as follows. Section 2 discusses the effect of atom loss and its detection using state-selective readout, and the delayed-erasure decoder developed in this work. In Section 3, we tailor and compare different SE techniques designed to handle loss with minimal overhead, which are necessary in circuits with high depth before logical measurements. Section 4 explores the optimal frequency of interleaving loss-detecting SE with transversal gates in deep logical Clifford circuits. In Section 5, we highlight the prevalent use of teleportation in many algorithmic subroutines and evaluate its effect with and without the delayed-erasure decoder on a toy model for the small-angle synthesis algorithm. Finally, we present our conclusions in Section 6.
2 Detecting and decoding delayed erasures
Figure 1: Loss errors in logical circuits. (a) Depiction of a logical algorithm with loss-detecting SE and gate teleportation. Physical qubit losses (red crosses) can generate correlated errors within and between logical qubits, complicating error correction. (b) Space-time diagram of a logical circuit, focusing on a measure qubit lifecycle during syndrome extraction. Physical qubits progress through time, undergoing initialization, gate operations, idling, and measurement. A loss event causes future gates to be cancelled, generating correlated errors between the qubits in the gate and flipping the corresponding stabilizers.
We consider general logical algorithms during which loss is periodically detected using state-selective readout (SSR), through loss-detecting SE and loss-detecting gate teleportation, as illustrated in Fig. 1(a). SSR corresponds to directly detecting atomic qubit states |$\ket{0}$, |$\ket{1}$, or atom loss. As it occurs as part of the qubit measurement, it can be realized through various methods in neutral atoms with minimal experimental overhead [43, 44, 11, 45, 46, 47, 24, 8]. A significant challenge in addressing loss errors in comparison to erasure errors is the uncertainty in the exact error location, which complicates the decoding procedure.
As a key quantity in predicting the performance of different loss detection and correction methods, we consider the qubit lifecycle, defined as the number of circuit locations where a given qubit can potentially be lost, starting at initialization and ending at measurement. For a typical realization of two-qubit gates [14], a neutral atom qubit which was expected to interact with a lost qubit experiences a single qubit error channel, rather than an entangling gate. Therefore, losing a qubit cancels all consecutive gates, leading to correlated Clifford errors within the circuit and causing stabilizers to flip (see Fig. 1(b)). In the context of logical algorithms, this can lead to correlated errors between logical qubits due to gate cancellation between logical qubits and propagation of errors during logical entangling gates (see Fig. 1(a)).
Figure 2: Delayed-erasure decoder. (a) Illustration of a qubit lifecycle and its usage in the delayed-erasure decoder. From initialization to measurement, each physical qubit can be lost at multiple possible time points, each occurring with a potentially different probability and corresponding syndrome. Upon detection of a qubit loss, the decoder accounts for these possibilities in order to improve the accuracy of the assigned correction. (b) Logical error rate for a logical memory as a function of the number of conventional SE rounds before logical measurement, here with distance $d=5$ and loss errors only with probability $p_{\text{loss}}=1\%$ per entangling gate. The delayed-erasure decoder (pink) substantially outperforms a decoder which does not account for loss information (black). Even without accurate loss information, the delayed-erasure decoder achieves comparable performance to a decoder with perfect loss information (gray, dashed).
While in principle these errors can be corrected by detecting the atom loss events and treating them as erasure errors, unlike the ideal erasure channels, where the location of the loss error is precisely known, in a natural QEC cycle, the loss event is detected when the SSR is performed and can correspond to a number of different potential loss locations (Fig. 2(a)), each resulting in a potentially different set of correlated errors. These loss detection events can be viewed as delayed erasure detections, and the information from SSR can be leveraged by the appropriate decoder to improve logical performance.
As a result, loss events require two adjustments to the decoder: first, when a qubit is lost, its associated stabilizer checks are no longer valid. Therefore, they need to be replaced with a so-called “supercheck”, a product of multiple stabilizer checks into a single check which is independent of the lost qubit [48, 49]. For example, consider a lost qubit $q_{4}$ that participates in neighboring stabilizers $S_{1}=Z_{1}Z_{2}Z_{3}Z_{4}$ and $S_{2}=Z_{4}Z_{5}Z_{6}Z_{7}$. The resulting supercheck operator is $S_{1}S_{2}=Z_{1}Z_{2}Z_{3}Z_{5}Z_{6}Z_{7}$, which is independent of the lost qubit $q_{4}$. Second, loss causes correlated errors, necessitating updates to the circuit error model to account for the likelihood of error propagation from the possible loss locations (see Supplementary Materials [50] Section S1 for further details).
We develop a delayed-erasure decoder which includes these two adjustments to effectively utilize imperfect information about the location of losses in time obtained from SSR. In practice, these adjustments can be paired with any decoding algorithm to augment its performance in the presence of loss. Here we primarily use the correlated most likely error (MLE) decoder from Ref. [12], as it can be applied to a broad class of codes and logical algorithms, and present results for the minimum weight perfect matching decoder in Supplementary Materials [50] Section S4.2.2 [51, 52]. Unlike previous approaches to decoding loss errors [21, 9, 22], our decoder automatically adjusts the error model based on the circuit and loss information, eliminating the need for hand-tuned models that may not easily generalize to complex logical algorithms or different SE methods.
Formalizing our approach, our goal is to automatically construct a decoding graph based on the observed loss, which captures how errors (hyperedges) trigger checks (vertices that compare consecutive stabilizer measurements in time). Ideally, we would solve the MLE decoding problem, which identifies the most likely configuration of Pauli and loss errors consistent with both the observed error syndromes and loss events (see Appendix Section BB.1). However, since a full MLE solution would require considering all combinations of loss locations due to the non-additive nature of loss-induced errors, this approach is computationally intractable for even modest system sizes and low loss rates.
Instead, we approximate the MLE solution by handling each loss event independently (see Appendix Section BB.2 for full details). For each loss event, we trace back the qubit’s lifecycle, accounting for all potential loss events and their associated probabilities. Each potential loss event corresponds to a loss circuit, in which certain gates are canceled due to the loss. Using Stim [53], a Clifford circuit simulator, we construct the hyperedges and their probabilities for each loss circuit. Next, we integrate all hyperedges within a qubit’s lifecycle, re-weighting them based on the probabilities of their corresponding potential loss locations. In cases of multiple losses, each loss event is calculated independently, and the results are combined to construct the final decoding graph, which serves as the input to the decoder. Empirically, this heuristic achieves performance comparable to methods that consider combinations of loss events, while significantly reducing computational overhead.
To benchmark the performance of our decoder, we focus on the well-studied case of a logical memory experiment on a surface code, which involves repeated SE rounds over time, while loss is detected only by SSR. In this setting, losses on measure qubits are detected every SE round, but losses on data qubits are not detected until the logical qubit is projectively measured at the end of the circuit. We perform simulations on a single surface code with code distance $d=5$ and circuit-level noise with an entangling gate loss error of $1\%$. Unless otherwise noted, this paper considers loss and Pauli errors occurring during entangling operations, as this represents a dominant source of error in many quantum computing architectures [4, 9, 14]. Details of the error model, along with results using alternative error models, can be found in Supplementary Materials Section S4.2.3 [50] and Table 1.
We begin by analyzing the case without any loss replacement, while increasing the number of rounds—and therefore, the lifecycle length. Figure 2(b) shows the logical error rate for three types of decoders: the MLE decoder, which does not account for loss information (black lines); the delayed-erasure decoder, which uses SSR information (pink lines); and an erasure decoder with perfect loss location information (gray dashed lines). Despite lacking accurate loss information, the delayed-erasure decoder improves the logical error rate by several orders of magnitude. Furthermore, even for deep circuits with long lifecycles, its performance is comparable to that of decoding with perfect loss information. As we will see later, for shallower circuits, the performance of the delayed-erasure decoder approaches that achieved with both perfect loss location information and full lost qubit replacement every round. This is because, in shallow circuits, the propagation of loss errors is nearly deterministic, significantly simplifying the decoding problem and allowing the delayed-erasure decoder to effectively handle these errors.
In contrast, when the circuit depth is large prior to logical measurement, the qubit lifecycles are extended and loss errors can occur at multiple points. These losses introduce correlated errors, depending on when and where they occur. While the delayed-erasure decoder performs relatively well in this setting compared to a decoder which does not utilize loss information and also compared to the decoder with perfect loss information, its effectiveness deteriorates when lifecycles become too long due to error accumulation. To address this, loss detection and qubit replacement must be incorporated into the algorithm to maintain performance. Therefore, in the next section we explore practical approaches to detect and replace lost atoms with minimal overhead.
3 Techniques for addressing qubit loss in deep circuits
To effectively handle loss errors in deep circuits prior to logical measurement, we now explore several SE methods that simultaneously manage loss and Pauli errors. These methods differ in their qubit overhead, gate operations, and loss detection capabilities. We focus on three primary approaches: (1) conventional SE, which resembles traditional circuit-based quantum computing (CBQC) with various modifications for delayed-erasure conversion, (2) teleportation-based SE, which resembles measurement-based quantum computing (MBQC) [54, 55] and (3) modified Steane SE [32]. We find that the underlying circuits and resulting performance of the methods are similar, but not identical: all approaches successfully remove qubit loss and have trade-offs in logical error rates and qubit overheads depending on the specific ratios between loss and Pauli errors. We begin by detailing each SE approach before analyzing numerical results in Section 3.4. In Section 3.5, we relate these performance differences to key metrics such as lifecycle length and number of entangling gates to predict performance based on the noise model for each SE method.
3.1 Modified conventional SE
The conventional SE method involves repeated stabilizer measurements using physical measure qubits. Data qubits are not directly measured or replaced when lost, and thus over time, logical performance can degrade in the presence of loss. We consider augmenting conventional SE by utilizing physical SWAP gates to detect losses on all qubits (SWAP SE), as proposed in Refs. [56, 42] and further explored in Refs. [24, 23]. At the end of each SE round, a SWAP gate and physical SWAP movement are performed between data and measure qubits. Conveniently, this approach does not require any additional entangling gates by applying gate cancellation identities (see Fig. 3(a)). If the data qubit is not lost, the SWAP operations cancel each other, such that the resulting measure qubit stores the stabilizer outcome. Conversely, if the data qubit was lost, the loss is directly identified through the SSR measurement and automatically replaced. Thus, using this method, each qubit cycles through both roles of data and measure qubits, ensuring a uniform lifecycle length of $\sim 8$ for all qubits in the bulk, even in deep circuits (see Supplementary Materials  [50] Section S2.1). However, a data qubit loss detection comes at the cost of losing a stabilizer outcome information, potentially impacting performance, as observed numerically in the next section.
Each SWAP operation incurs a cost due to qubit movement, making it crucial to evaluate different SWAP periods to balance performance and experimental complexity. The SWAP period defines how often loss-detecting SE rounds are interspersed with conventional SE rounds, occurring at a fixed interval. We identify a trade-off between the lifecycle lengths of data and measurement qubits for different SWAP periods, demonstrating that a period of 2 can be competitive with a period of 1 (see Supplementary Materials [50] Figs. S2, S3). A period of 2 reduces experimental complexity while maintaining a similar average lifecycle length.
3.2 Teleportation-based SE
Teleportation-based SE is an alternative QEC approach which utilizes regular teleportation between logical qubit layers, inherently enabling loss detection upon logical measurement [54, 55, 22, 16] (see Fig. 3(b)). While typically associated with measurement-based quantum computation (MBQC) architectures, the teleportation-based SE can be used in a circuit-based framework where gates are directly applied on qubits. We realize the XZZX cluster state [30], as described in detail in the Supplementary Materials [50] Section S2.5. In this framework, each qubit is entangled with four neighboring qubits, which, under our error model considering noise on entangling gates only, results in short lifecycles of length four. Among the SSR-based approaches considered, this SE method enables the shortest lifecycles with no experimental overhead apart from SSR, though it requires additional qubits. These features make it well-suited for neutral atom architectures and particularly effective in loss-dominant regimes, though its advantages diminish when the loss fraction is small, as detailed in Section 3.4.
3.3 Modified Steane SE
Steane SE uses transversal $CX$ gates applied to fault-tolerant logical measure qubits to extract syndromes and detect errors [32]. Transversal gates, which apply the same operation across corresponding physical qubits in each code block, ensure that errors propagate predictably from logical data qubits to logical measure qubits. If the measure qubits are prepared fault-tolerantly with $d$ SE rounds, one round of Steane SE suffices to accurately capture the syndromes in a given basis [32]. However, conventional Steane SE lacks a native mechanism for detecting loss, as losses on data qubits are not detected by the logical measure qubits and remain hidden until the end. We modify the conventional Steane SE approach by incorporating logical SWAP operations (SWAP gate and SWAP movement), as illustrated in Fig. 3(c), thereby enabling loss detection through SSR at each logical measurement. This approach leverages the teleportation of logical information and resembles Knill SE [57].
The modified Steane SE can be thought of as an interpolation between modified conventional SE and teleportation-based SE, similar to the ideas in Ref. [58]. In the limit of multiple logical measure qubits, each prepared with one SE round, the modified Steane SE becomes equivalent to teleportation-based SE (see Supplementary Materials [50] Section S2.6). Conversely, multiple rounds of SE can be used to prepare a higher-quality logical measure qubit, providing flexibility to adapt to various error models and resource constraints. Steane SE also supports pre-selection of measure qubit blocks based on quality, potentially increasing fidelity without significant overhead [59, 60, 61, 62]. To provide a comprehensive perspective on SE techniques, we include Steane SE here despite not numerically analyzing it in the main text. Since modified Steane SE interpolates between modified conventional SE and teleportation-based SE, we expect its performance to qualitatively reflect this interpolation. Thus, we highlight its conceptual advantages as a potential direction for future studies.
Figure 3: Loss-detecting SE methods. (a) Modified SWAP SE method for detecting and correcting loss errors using physical SWAP operations between data and measure qubits, without gate overhead. (b) Teleportation-based SE method, utilizing layers of teleportation and stabilizer checks in alternating bases. The circuit is repeated $d$ times. (c) Steane SE method with logical SWAP, employing logical fault-tolerant measurement qubits and teleportation. (d) Logical error rates as a function of SE rounds using the delayed-erasure decoder, with a code distance of 7 and a physical error rate $1\%$. Comparable performance is observed across SE methods in regimes with short lifecycles. (e) Error thresholds as a function of the loss fraction for different SE methods, showcasing the improvement with loss for all methods. (f) Effective distance as a function of loss fraction for different SE methods, considering distance $d=7$. (g) Space-time overhead as a function of the loss fraction, describing the required distance to achieve logical error rates of $10^{-12}$ given a physical error rate of $0.5\%$. The legend for (d-g) appears in (e).
3.4 Comparison of SE methods
We now compare the performance of these approaches in the context of a surface code memory, using circuit-level simulations with varying loss fractions $L={p_{\text{loss}}}/({p_{\text{loss}}+p_{\text{Pauli}}})$, where $p_{\text{loss}}$ and $p_{\text{Pauli}}$ are physical loss and Pauli error probabilities, respectively. We identify the optimal SE approach at different loss fractions and Pauli error biases (see Supplementary Materials [50] Section S4 for detailed results with bias).
We also evaluate a Free SE method, which detects and replaces lost qubits at different intervals: after every gate (period 0.25) or every $K$ rounds (period $K$). Free detection at period 0.25 eliminates loss propagation, ensuring that the exact loss locations are known, effectively realizing an erasure channel. Furthermore, we compare our results to a Free SE method with period 1, including an additional feature: the exact loss locations are known, greatly simplifying the decoding problem. A possible implementation of these SE approaches involves mid-circuit erasure conversion and atom replacement [17].
Figure 3(d) compares the logical error rates of conventional SE without loss detection, approaches with loss detection—such as SWAP SE, teleportation-based SE, and Free SE—and Free SE with perfect loss information, in a single logical memory experiment at an experimentally-motivated loss fraction of 0.5. For a small number of SE rounds ($\leq 10$), all protocols behave similarly. However, as the number of SE rounds increases, loss detection and the replacement of lost qubits substantially enhance performance, providing a convenient approach to achieving high circuit depths without additional experimental gate overhead.
Figure 3(e) shows the threshold of each SE method as a function of the loss fraction. We determine the threshold by calculating the logical error rate for different code distances $d$, using a noiseless initialization, followed by $d-1$ rounds of noisy stabilizer measurements and a final noiseless transversal measurement. We find that the thresholds of all methods improve with increasing loss fraction. Notably, methods with shorter lifecycles, such as teleportation-based SE, benefit more from increasing loss fraction. In the next Section 3.5, we provide further insight into the thresholds of each SE method by linking it to simple characterizations such as qubit lifecycles.
Figure 3(f) presents the effective code distance of each SE method as a function of the loss fraction, for distance $d=7$. The effective distance $d_{e}$ refers to the number of errors required to cause a logical failure, as the logical error rate scales as $(p/p_{\text{th}})^{d_{e}}$ far below the threshold $p_{\text{th}}$, where $p$ is the physical error rate. For Pauli noise, $d_{e}=(d+1)/2$, and for erasure noise $d_{e}=d$. To determine $d_{e}$ for each loss fraction, we fit the logical error rate data far below the threshold to the function $\alpha p^{\beta}$, where $\alpha$ and $\beta$ are fitting parameters (see Supplementary Materials [50] Fig. S13). As expected, all SE methods, while utilizing the delayed-erasure decoder, experience increased effective distance with an increased loss fraction. For loss errors only, teleportation-based SE and Free SE methods (periods 0.25 and 1) achieve the optimal effective distance of $d_{e}\approx d$, while SWAP SE achieves $d_{e}\approx d-1$, likely due to longer lifecycles. Additional threshold and effective distance results under an alternative error model from Ref. [22] are presented in Table 1.
To account for the qubit overhead required in different SE methods, we evaluate the space-time overhead for performing $d$ SE rounds of each method in Figure 3(g). For a given physical error rate of $0.5\%$, we determine the required code distance to achieve a logical error rate of $10^{-12}$ by fitting the logical error as a function of distance for $d=3,5,7,9$. We then present the space-time overhead of each approach as a function of the loss fraction, highlighting the overhead reduction as the loss fraction increases for all SE approaches. Our results highlight that although teleportation-based SE requires more qubits, its better error suppression when the loss fraction is high leads to a more favorable space-time volume compared to SWAP SE (see space-time overhead for each SE method in Table 1).
An important regime is where the loss fraction is $\approx 0.5$, which is comparable to the ratio observed in recent neutral atom experiments [14, 10]. In this regime, the thresholds of the different methods range from 1.5% to 2.5%, and all SE methods achieve similar effective distances. However, these methods can vary significantly in experimental complexity. In particular, erasure conversion requires directly detecting and replacing the qubit erasure without projectively measuring it, unlike with SSR [17]. Clearly, all methods provide benefits when the loss is correctly managed and detected, and experimental considerations can illuminate which may work best in practice.
Another important aspect of neutral atom qubits is their intrinsic bias, as $Z$-type Pauli errors are much more common than $X$-type Pauli errors [14]. In Supplementary Materials Section S4.2.2 [50], we investigate the interplay between biased errors and delayed erasure errors across multiple SE methods, employing the XZZX surface code [28] and the XZZX cluster state [30] for teleportation-based SE. Our analysis considers scenarios both with and without bias-preserving gates, as well as the presence of biased erasure noise. Bias-preserving gates, such as native $CX$ gates, preserve the bias of the errors, in contrast to $CX$ gates decomposed into $CZ$ and $H$ gates. Biased erasure noise occurs when a qubit exits the computational subspace exclusively from the state |$\ket{1}$. As a result, the replaced qubit follows a biased error channel instead of a depolarizing error channel. Additionally, we investigate scenarios where Pauli errors are biased, examining the two-dimensional space defined by the bias ratio and the loss fraction. Thresholds are calculated at each point in this space for various SE methods (see Figures S16 and S17 in the Supplementary Materials [50]). A key observation emerges across all SE methods: increasing the loss fraction has a significantly greater impact on thresholds than increasing the bias ratio, even when using bias-preserving gates and biased-erasure. This difference in impact between the loss fraction and the bias ratio becomes even more pronounced in the absence of bias-preserving gates.
Figure 4: Linking thresholds to key metrics. (a) Thresholds for different SE methods for different loss and Pauli error rates. The curves are linear fits to numerical finite-size data, with the region below each curve representing the correctable region. (b) Thresholds as a function of lifecycle length for various SE methods, in the loss error only limit.
3.5 Predicting performance by error counting
We now present a unified model for analyzing SE methods by linking performance to simple characterizations such as lifecycle length and the number of entangling gates. Figure 4(a) shows the thresholds in the parameter space of loss and Pauli error rates for various SE methods. Thresholds are plotted in terms of the loss error rate and Pauli error rate, with the region below each curve representing the range of correctable errors. We numerically determine a good fit to a linear model (solid lines) based on finite-size data, given by $p_{\text{loss}}=p_{\text{loss,th}}-(p_{\text{loss,th}}/p_{\text{Pauli,th}})\cdot p_{\text{Pauli}}$, where $p_{\text{loss,th}}$ and $p_{\text{Pauli,th}}$ are the respective thresholds for loss and Pauli errors only. The linear behavior suggests that the threshold depends on the loss fraction $L$ according to the relationship $p_{\text{threshold}}=p_{\text{loss,th}}p_{\text{Pauli,th}}/(L(p_{\text{Pauli,th}}-p_{\text{loss,th}})+p_{\text{loss,th}})$, and is qualitatively similar to the behavior observed in Refs. [49, 63, 48, 23].
The curve intersections with the axes provide key insights. The y-axis intersection, representing the threshold in the absence of Pauli errors, is related to the lifecycle length. For the Free SE period 1 with perfect loss information, the threshold is $p_{\text{loss,th}}\approx 7.2\%$. Increasing the lifecycle reduces the threshold, which decays with the lifecycle length, as shown in Fig. 4(b) (see Supplementary Materials [50] Section S2.7 for further details). Notably, the SWAP SE, with an average lifecycle length of 8, has a lower threshold of $p_{\text{loss,th}}\approx 2.3\%$, deviating from the heuristic. This lower performance is attributed to additional factors unique to SWAP SE, such as the inability to detect measurement qubit loss when data qubits are detected as lost and the loss of stabilizer information, as discussed previously in Section 3.1.
The x-axis intersection, representing Pauli error thresholds, correlates with the number of entangling gates, which impacts the number of Pauli errors in the final state. For example, over $d$ SE rounds, teleportation-based SE uses 1.5×1.5$\times$1.5 × more entangling gates than SWAP SE and Free SE approaches, resulting in a proportional decrease in the Pauli error threshold, as similarly observed in Ref. [16].
4 The effect of Loss Errors in deep Logical Algorithms design
Figure 5: Deep logical circuits with qubit loss. (a) A deep Clifford logical algorithm consisting of random logical single-qubit and transversal $CX$ gate layers, with periodic SE rounds at varying frequencies. (b, c) Circuit-level simulation results showing the logical error rate as a function of the number of SE rounds per transversal gate layer, for different loss fractions. ($p=1\%$, $d=5$, 24 layers). The SWAP SE method (c) effectively mitigates loss errors, restoring the optimal SE frequency observed in Pauli-dominated scenarios. In contrast, conventional SE (b) exhibits varying error correction regimes, where loss can either improve or degrade performance depending on the lifecycle length.
Loss errors in logical algorithms can have significantly different effects compared to standard memory benchmarks. We now study the effects of loss errors on the QEC design of multi-qubit deep logical algorithms. Specifically, we analyze how physical loss errors influence the optimal frequency of SE rounds between transversal gates. Our analysis reveals that SWAP-based SE achieves comparable optimal SE frequencies in both the presence and absence of loss.
We focus on the effects of loss in random Clifford logical algorithms (Fig. 5(a)), where multiple logical qubits interact through transversal gates interspersed with periodic SE rounds at tunable frequencies. Circuit-level simulations are performed for logical circuits comprising 24 layers of logical $CX$ and single-qubit logical gates ($X,Y,Z$), with SE rounds applied at varying intervals, as in Ref. [12]. Two scenarios are analyzed: conventional SE without explicit loss detection, which detects only measurement qubit loss in each SE round (Fig. 5(b)), and SWAP-based SE (Fig. 5(c)). The x-axis represents the number of SE rounds per gate layer, ranging from no SE ($n_{r}=1/24$) to multiple SE rounds after every gate layer ($n_{r}=3$).
In the absence of SWAP operations (but still leveraging SSR detection and delayed-erasure decoding), we observe several distinct error correction regimes. For a small number of SE rounds per gate, loss improves the logical error rate due to short qubit lifecycles. However, as the number of SE rounds per gate increases, performance degrades due to the cumulative effects of loss over longer lifecycles. By contrast, incorporating SWAP-based SE, loss consistently enhances error correction performance. Loss detection via SWAP stabilizer checks restores the previous optimal SE rounds per $CX$ observed in Ref. [12]. Consequently, when employing SWAP SE and SSR, the presence of loss does not alter the heuristic conclusions made for logical algorithms dominated by Pauli errors and can, in fact, improve overall performance.
5 Native loss detection from logical teleportation
Figure 6: Native loss detection in logical algorithms. (a) Logical qubit teleportation during gate teleportation, shortening lifecycles and naturally detecting loss without loss-detecting SE rounds. (b) SE rounds per logical qubit before logical measurement for various key algorithmic subroutines. Each subroutine uses frequent gate teleportation, keeping lifecycles short and detecting loss without additional experimental overhead.
While we have techniques for handling algorithmic structures with long qubit lifecycles, we now observe that in many cases, qubit lifecycles are short in realistic logical algorithms. In particular, teleportation is a powerful technique for loss detection and correction that avoids additional overhead, and it naturally emerges within logical algorithms through gate teleportation (see Fig. 6(a)). Specifically, the SWAP-teleported gates exchange the logical data qubit with a teleported logical qubit that implements the desired gate, detecting loss using SSR and terminating the qubit lifecycles without the need for additional SE rounds.
This approach is further enhanced by correlated decoding and algorithmic fault tolerance, which capitalize on the use of transversal gates and teleportation in universal quantum computation [12, 13]. These ensure that lifecycles are inherently short in many logical algorithms, simplifying loss management while enabling regular qubit replacement.
A key observation is that a wide range of known logical subroutines naturally employ teleportation, inherently keeping qubit lifecycles short. Fig. 6(b) illustrates the average number of SE rounds per physical qubit, from initialization to measurement, across various essential subroutines. The results show that most algorithms have relatively brief lifecycles. Detailed descriptions of these algorithms are provided in the Supplementary Materials [50] Section S3. For this analysis, we conservatively assume one SE round per logical operation. However, as shown in Fig. 5(c), this number can be further reduced. Additionally, Fig. 3(d) and Supplementary Figure S11 [50] reveal that, at an experimentally-motivated loss fraction of 0.5, for fewer than 10 SE rounds, performance remains nearly unchanged without any active loss correction added to the SE rounds, relying solely on SSR and delayed-erasure decoding. These findings suggest that while memory benchmarks provide useful performance metrics, transversal logical algorithms are inherently well-equipped to manage loss, with minimal to no loss detecting SE. Leveraging SSR and delayed-erasure decoding can significantly enhance the performance of the logical algorithm without incurring additional experimental complexity.
Figure 7: Deep logical algorithms with teleported gates. (a) We study a deep circuit consisting of teleported $X$ and $Z$ logical gates, similar in structure to small-angle synthesis algorithms [35, 36]. (b) Circuit-level simulation results for 11 layers of random teleported $Z$ and $X$ logical gates, with all logical qubits initialized in the presence of noise (physical error rate $p=1\%$ and $d=7$). The delayed-erasure decoder (pink) substantially outperforms an MLE decoder that does not account for loss (black) and improves with increasing loss fraction. It also matches the performance of a decoder with perfect loss information (gray, dashed), and for loss fractions less than 1, approaches that of the erasure channel (khaki, dashed). The erasure channel performance is achieved using Free SE with period 0.25, corresponding to loss detection and qubit replacement after every gate, thus providing a lower bound.
As a proof of concept, we consider the teleportation-based logical circuit shown in Fig. 7(a), which mirrors the structure of the small-angle synthesis algorithm [35, 36]. This algorithm constructs small-angle rotations through sequences of $H$ and $T$ gates, with the latter often realized via teleportation. Consequently, loss detection and correction are integrated into the process of executing the logical algorithm. The logical error rates for the circuit in Fig. 7(a) are plotted in Fig. 7(b) as a function of the loss fraction using various decoders: a delayed-erasure decoder leveraging SSR, a regular MLE decoder, and an MLE decoder with perfect loss location information. The delayed-erasure decoder demonstrates significant performance improvements, with algorithmic logical error rates rapidly decreasing as the loss fraction increases, achieving the performance of the decoder with perfect loss information. Furthermore, for loss fractions less than 1, it achieves the lower bound set by the erasure channel—corresponding to loss detection and qubit replacement after every gate. Notably, this circuit is similar to the teleportation-based SE method investigated here in the memory setting, both utilizing SSR and replacing atoms frequently through teleportation without additional overhead. These results highlight the advantages of exploiting native loss detection within teleportation gadgets in logical circuits while also utilizing loss information to improve decoding. Finally, while our simulations focus on Clifford gate teleportation, they are expected to extend to non-Clifford gates, as the underlying loss detection principles remain consistent.
6 Conclusion
These considerations provide a framework for managing atom loss in neutral atom QEC, and explore the role of loss detection and bias in improving logical algorithm performance. Central to our approach is the use of the delayed-erasure decoder, which leverages loss information to approximate an MLE decoding solution, substantially improving the logical error rate.
Using our decoder, we examine how the algorithmic structure influences the optimal strategy for detecting and correcting loss. For high-depth circuits involving a large number of gate layers prior to logical measurement, the performance depends on both loss decoding and frequent loss detection and replacement. Using a surface code logical memory, we explore different SE methods under various loss fractions. Our findings indicate that, with appropriate decoding strategies, all methods enhance performance with increasing loss fraction. Notably, teleportation-based SE is a promising candidate for neutral atom quantum computing, as it leverages SSR to achieve comparably high thresholds for high loss fractions, albeit with an additional space overhead. Determining the optimal SE method for a given system will depend on experimental validation and the specific noise characteristics of the hardware.
By applying these results to multi-qubit deep logical algorithms, several key insights emerge. Atom loss errors, when managed using loss detection operations and SSR detection, are fully compatible with correlated decoding, allowing for an optimal number of approximately four logical operations per SE round using the delayed-erasure MLE decoder. This framework ensures effective error correction even in the presence of high loss rates.
By considering key subroutines involving extensive logical teleportation, we find that the use of transversal gates and correlated decoding keeps the number of SE rounds sufficiently small, such that loss is natively detected and managed by the gate teleportation intrinsic to universal processing. Examining a toy model of a small-angle synthesis algorithm, we observe that loss errors, decoded using our delayed erasure decoder, significantly enhance performance compared to Pauli channels solely through logical teleportation. As such, while the logical memory benchmark provides valuable insights into various strategies for loss management and detection to enhance performance, behavior can radically change when realizing logical algorithms.
This work opens up several new research directions. A major priority is reducing the decoding runtime, critical for scaling to more complex and larger algorithms. While the present work mostly employs an MLE inner decoder, its limited scalability for large-scale logical algorithms underscores the need for further refinement, possibly through alternative strategies, such as matching decoders [64, 65, 9] or machine learning-based decoders [66, 9].
An intriguing avenue for future exploration involves optimizing loss detection in specific algorithmic subroutines, tailoring the loss detection frequency and SE method to align with the algorithmic gadget and system characteristics. Additionally, the delayed-erasure decoder can be naturally applied to the decoding of transversal non-Clifford circuits, as it is inherently compatible with and effective for correcting Clifford errors propagating through the circuit. Finally, our decoding techniques are broadly applicable to a wide range of QEC codes and logical algorithms, making them well-suited for further exploration of loss decoding in high-rate low-density parity-check (qLDPC) codes  [67, 68, 69, 70, 71, 72, 73, 74, 75].
Note added. Since the completion of our work, we became aware of several related studies examining the effect of qubit loss on a single logical memory: teleportation-SE [22], loss detection units [23], and SWAP-SE [76]. While addressing similar challenges, these works differ from ours in focus and methodology. Notably, improved thresholds have been recently reported in updated versions of Refs.[22, 76], based on the use of decoding strategies similar to the delayed erasure decoder described in this work. These contributions also utilize error models slightly different from those presented here. For completeness, our updated Appendix A includes results obtained using the error models utilized in Refs.  [22, 76].
Acknowledgments
We thank Sasha Geim, Alex Kubica, Jeff Thompson, Liang Jiang, Isaac Chuang, Nazli Ugur Koyluoglu, Luke Stewart, Marcin Kalinowski, Edita Bytyqi, David C. Spierings, Katie Chang, Casey Duckering, Shengtao Wang, Chen Zhao for insightful discussions and Susanne Yelin for her guidance and valuable discussions. G.B. acknowledges support from the MIT Patrons of Physics Fellows Society. D.B. acknowledges support from the NSF GRFP (grant DGE1745303) and the Fannie and John Hertz Foundation. We acknowledge financial support from IARPA and the Army Research Office, under the Entangled Logical Qubits program (Cooperative Agreement Number W911NF-23-2-0219), the DARPA IMPAQT program (HR0011-23-3-0012), and the DARPA MeasQuIT program (HR0011-24-9-0359), QuEra Computing Inc (award number A57912), the Center for Ultracold Atoms (an NSF Frontier Center), and the National Science Foundation (award number PHY-2012023).
Appendix A SE Methods Lifecycles and Space-time Overheads
Category Conventional SWAP period 1 Free period 2 Teleportation-based Free period 1 Free period 1 + exact loss info Free period 0.25 + exact loss info
Average lifecycle length of data qubits $4d$ 8 8 4 4 1 (detection) / 4 (replacement) 1 (detection) / 1 (replacement)
Average lifecycle length of measure qubits 4 8 4 4 4 1 (detection) / 4 (replacement) 1 (detection) / 1 (replacement)
Space overhead $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 3d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$ $\sim 2d^{2}-1$
Time overhead $\sim 4d$ $\sim 4d$ $\sim 4d$ $\sim 6d$ $\sim 4d-1$ $\sim 4d-1$ $\sim 4d-1$
Extra experimental requirements None SSR Erasure conversion [17, 10], detection and replacement every 2 SE rounds SSR Erasure conversion [17, 10], detection and replacement every SE round Erasure conversion [17, 10], detection every gate, and replacement every SE round Erasure conversion [17, 10], detection and replacement every gate
Threshold for $L=1$ [%] (Channel 1) NA $2.27\pm 0.02$ $3.98\pm 0.06$ $4.78\pm 0.01$ $5.07\pm 0.02$ $7.16\pm 0.01$ $9.51\pm 0.04$
Effective distance for $d=7$ (Channel 1) NA $6.41\pm 0.27$ $6.77\pm 0.2$ $7.19\pm 0.45$ $6.83\pm 0.37$ $6.97\pm 0.24$ $7.46\pm 0.36$
Threshold for $L=1$ [%] (Channel 2) NA $1.94\pm 0.01$ $2.75\pm 0.07$ $3.5\pm 0.05$ $3.76\pm 0.01$ $5.12\pm 0.01$ $6.5\pm 0.01$
Effective distance for $d=7$ (Channel 2) NA $6.13\pm 0.26$ $5.5\pm 0.23$ $6.5\pm 0.21$ $6.32\pm 0.87$ $6.45\pm 0.47$ $7.47\pm 0.49$
Table 1: Comparison of SE methods under various error models. Each column represents a method, and each row describes a property such as average qubit lifecycles, thresholds, effective distance and overheads. The final column estimates performance under very frequent detection and replacement, providing the erasure channel bound. Channel definitions for $L=1$ are as follows. Channel 1: $p^{\prime}\{L\otimes I,I\otimes L\}$. Channel 2: $\frac{p^{\prime}}{2}\{L\otimes Z,Z\otimes L,L\otimes I,I\otimes L\}$. $p^{\prime}=1-\sqrt{1-p}\sim p/2$ where $p$ is the entangling gate physical error rate.
Table  1 summarizes the average lifecycles, error thresholds, and space-time overheads for each SE method.
Appendix B Delayed-Erasure Decoder
In this section, we will describe our approach for decoding in the presence of loss. We first describe the exact MLE solution. Then, we describe the approximated MLE decoding which we use in the main text.
B.1 Exact MLE Decoding in the Presence of Loss
In this section, we provide a formulation of the most-likely-error (MLE) decoding problem in the presence of loss errors. Loss acts in a much more complex fashion compared to normal Pauli errors, and therefore require more care in terms of their formulation. The results also depend on the details of the error model and measurement model, introducing further complications. We will attempt to describe the full formulation in this section, which will then guide the choice of heuristic methods, primarily based on reweighting edges in the decoding graph, that will approximate the MLE problem.
We formulate the problem by generalizing the discussion in Ref. [12]. We can associate a binary variable $E_{i}$ for each Pauli error, and similarly $L_{i}$ for each loss event. We assume that the loss and Pauli errors are applied independently (i.e. the channel does not simultaneously apply two types of errors, even though that could be the case when doing a CZ gate, where there is a correlation between loss on one qubit and a Pauli error on the other). When running the circuit, we obtain information about errors and loss through qubit measurements. This information comes in two flavors: First, there are flags $F_{j}$, which indicate a qubit was lost when we tried to measure it. Second, there are detectors $D_{j}$, which are products of stabilizer measurement results that should be deterministic in the absence of errors. A non-trivial aspect of loss decoding is that the detectors must be determined based on the flags, since some qubit measurement results are labeled as “invalid” instead of 0 or 1.
The most likely error problem seeks to find the assignment of Pauli errors $\vec{E}$ and loss errors $\vec{L}$ that maximizes the following quantity:
$P(\vec{E},\vec{L}|\vec{F},\vec{D}),$ (1)
Namely, given the observed flags, and the resulting detectors that we define conditional on seeing those flags, what is the most likely error.
We can rewrite this expression using Bayes’ rule
$P(\vec{E},\vec{L}|\vec{F},\vec{D})$ $=\frac{P(\vec{E},\vec{L},\vec{F},\vec{D})}{P(\vec{F},\vec{D})}=\frac{P(\vec{F},\vec{D}|\vec{E},\vec{L})P(\vec{E},\vec{L})}{P(\vec{F},\vec{D})}.$ (2)
Since we are given a fixed observation $\vec{F}$, $\vec{D}$, the denominator is fixed. We therefore seek to maximize the product of the probability of the error configuration $P(\vec{E},\vec{L})$ and the conditional probability of observing these syndromes. In the presence of loss, an important distinction from the usual case is that the latter factor is no longer always 0 or 1.
The error probability $P(\vec{E},\vec{L})$ can be readily written down based on the log likelihood ratios of individual events. One potential subtlety that one may need to worry about is that for a given qubit at a given lifecycle, only one loss event can occur. Thus, if we have a sequence of gates with loss probability $p_{1}$, $p_{2}$, …, then the loss event probabilities may need to be written as $p_{1}$, $(1-p_{1})p_{2}$, $(1-p_{1})(1-p_{2})p_{3}$,…, although the deviations may be higher-order in practice.
The more challenging factor to evaluate is $P(\vec{F},\vec{D}|\vec{E},\vec{L})$. First, we need to specify how the new detectors are chosen given a heralded loss pattern. In Ref. [42], they compare the stabilizer value against the next available syndrome measurement result at the same site. This seems like a reasonable strategy, since this bit of information is really not accessible until the next time we measure this stabilizer, and is not revealed by any other stabilizers. Note that strictly speaking, if we solve the full MLE problem, the precise choice of detector shouldn’t matter due to stabilizer equivalences, so long as it forms a complete basis. However, it will matter if we are using heuristic decoders. After forming the detectors, we need to evaluate whether a given loss and error pattern is consistent with the observed loss flags and detectors. The loss flag check is straightforward, and can be imposed as a constraint that exactly one of the loss locations corresponding to a given qubit are activated. The detector is more subtle, because now some of the way they are triggered becomes probabilistic. A brute force way to check would be as follows: For a given loss pattern, the circuit is now known. We can forward propagate all error events through the new circuit, and obtain measurement results, which will then tell us the probability distribution and allow us to calculate the conditional probability. Unfortunately, this seems rather costly to perform and may not be as easily phrased as a simple optimization problem with a linear optimization target, so we need to think a bit more about better ways to evaluate this probability.
Finally, we comment on the benefits and downsides of heuristic edge re-weighting techniques. For example, in Ref. [42], given a leakage event, they assign an $X$ and $Z$ error on the $i$th location with probability $i/2n$ and propagate this error forward to modify the edge weights. Note that the detailed error event is related to their leakage model, where all qubits that interact with a leaked qubit experience large depolarizing noise. This is good to leading order, since it captures the fact that qubits that interact with the measure qubit later are more likely to have experienced leakage. However, it seems to still miss one potentially important aspect: If a qubit is leaked at some point, then it is leaked for the rest, so the error events are actually correlated. This correlation can in principle be leveraged to more accurately deduce the error that occurred.
Given these considerations, the main goal is thus to find good heuristic methods that approximate MLE. In particular, some better intuition about the types of error events that are incorrectly classified for a given weighting could help inform modifications that make it closer to MLE. One could also either try different heuristics and see what gives the best logical error rate (focusing perhaps on “hard” instances to make the search more efficient), or implement MLE so that it also becomes clear how big the optimality gap to MLE is for a given heuristic (this also allows one to more clearly focus on concrete instances of heuristic decoding failures). With the possible exception of neural network decoders that automatically handle both aspects (see Google’s paper, where there are actually also simulations with heralded leakage [5]), it seems overall that closing the gap to MLE may make a bigger difference than the MLE-MLD gap.
B.2 Approximate MLE Decoding
In this section, we describe our approach for constructing a decoder to handle losses by approximating the MLE decoding introduced in the previous section.
We utilize the concept of qubit lifecycles: the number of potential locations where a physical qubit may be lost—due to single-qubit gates, multi-qubit gates, initialization, idling, measurement, or other processes. Given a heralded loss upon measurement, the decoder identifies a loss occurring in one of these locations, each characterized by its specific probability of loss.
Consider a scenario where, at the end of qubit $i$’s lifecycle $C_{i}$, a loss is detected, with $L_{ij}$ representing potential loss events, each associated with a probability $p_{ij}$. For each potential loss event $L_{ij}$, we simulate the effect of the loss by constructing a loss circuit. This involves removing subsequent gates acting on the lost qubit and modeling the measurement of the lost qubit as random. Using Stim’s gauge detector infrastructure [53], we generate the decoding graph for each specific loss event, denoted as $\text{DEM}_{ij}$.
The decoding graph is a collection of hyperedges and their associated probabilities. Each error mechanism corresponds to a hyperedge that connects the detectors it triggers. Detectors are defined as the products of stabilizer measurements and are expected to be $+1$ in the absence of errors.
Each decoding graph $\text{DEM}_{ij}$ is a collection of error probabilities $p_{n}$ and corresponding detector configurations $D_{n}$:
$\text{DEM}_{ij}=\{(D_{n},p_{n})\},$ (3)
where $D_{n}=[D_{n}(0),D_{n}(1),\cdots]$ represents the list of detectors triggered by the error event. Repeating this process for all potential loss events $L_{ij}$ within a lifecycle $C_{i}$ results in a set of decoding graphs $\text{DEM}_{ij}$ and their associated probabilities $p_{ij}$.
To compute the final decoding graph for a lifecycle $C_{i}$, we sum over all decoding graphs for the potential loss events:
$\text{DEM}_{i}=\sum_{j}p_{ij}\cdot\text{DEM}_{ij}.$ (4)
This summation is straightforward because only one of these events can occur.
When multiple qubit losses occur in different lifecycles, loss errors can cause correlated Clifford errors. Combining syndromes from distinct lifecycles may interfere non-linearly, resulting in syndromes that cannot be derived from independent consideration of loss events. While the optimal decoder would consider all combinations of potential loss locations for all qubits, this approach scales poorly with multiple lossy lifecycles and potential loss events, making it impractical for small code distances ($d<7$).
To address this, we explore a decoder that evaluates each lifecycle independently and averages over them. Additionally, we introduce the option to include a single combination of losses across different lifecycles. Specifically, we consider the first potential loss locations $L_{i1}$ for all lossy lifecycles $L_{i}$ and generate a corresponding decoding graph, $\text{DEM}_{\text{first comb}}$.
The decoding graphs for individual potential loss locations $L_{ij}$ are generated in a pre-processing step, independent of any specific error model or probabilities of Pauli and loss errors. During real-time decoding, these graphs are summed using the specific error model probabilities $p_{ij}$. Note that we separately account for loss errors and Pauli errors. Lossless circuits (without losses) are used to efficiently generate the decoding graph for Pauli errors, denoted $\text{DEM}_{\text{Pauli}}$.
For a specific shot with a heralded loss pattern using SSR and a given error model, the final decoding graph is computed by summing: 1. Lossy lifecycle decoding graphs $\text{DEM}_{i}$, 2. The Pauli decoding graph $\text{DEM}_{\text{Pauli}}$, and 3. The first potential loss combination decoding graph $\text{DEM}_{\text{first comb}}$.
The final decoding graph is given by:
$\text{DEM}_{\text{final}}=\sum_{i}\text{DEM}_{i}+\text{DEM}_{\text{Pauli}}+\omega\cdot\text{DEM}_{\text{first comb}},$ (5)
where $\omega$ is a combination weight determined based on the analysis in Appendix Section B.2.B.2.1.
The summation of probabilities is calculated using the following equation:
$\sum_{i=1}^{l}p_{i}\prod_{j\neq i}(1-p_{j})+\sum_{i=1}^{l}\sum_{j=i+1}^{l}\sum_{k=j+1}^{l}p_{i}p_{j}p_{k}\prod_{m\neq i,j,k}(1-p_{m}).$ (6)
This ensures accuracy up to $O(p^{3})$.
Finally, given the final decoding graph, any known decoder can be used to decode the data for a specific shot, such as MWPM, MLE, or other methods.
B.2.1 Adjusting Loss Combination Weight
In this section, we explore the influence of the weight of the combination, $\omega$, on the efficacy of the delayed erasure decoder, which considers both independent error events and the first combination of potential loss events. As described earlier, the decoder’s decision-making process involves summing different decoding graphs (see Eq. 5). To incorporate the impact of individual loss events, the first loss combination event is scaled by $\omega$ along with the standard Pauli error decoding graph.
To benchmark the performance of our decoder and determine the optimal value of $\omega$, we performed circuit-level simulations on a memory logical qubit with multiple rounds of conventional syndrome extraction (SE). We varied $\omega$ between 0 and 1. For $\omega=0$, each lifecycle’s decoding graph is considered independently, without including an additional decoding graph for the first combination.
The results for the conventional SE method are presented in Fig. 8. Non-zero values of $\omega$ degrade the decoder’s performance. However, it is important to note that these results pertain to the conventional SE method, and different SE methods employing distinct gate sets may behave differently. For simplicity, the numerical simulations presented in this paper assume $\omega=0$.
Figure 8: Logical memory simulations with $d$ rounds of the conventional SE method. Variation of the logical error rate as a function of the combination weight $\omega$ for distance 9, across different physical error rates ($p=0.01$ and $p=0.03$). The results indicate that nonzero values of $\omega$ deteriorate the decoder’s performance.